OpenStudy (halox156):

What is the simplified form of X^4 - 81/x + 3?

OpenStudy (anonymous):

This is actually three different and unrelated questions. Let's concentrate on No. 1. The key in simplifying this is to recognize a characteristic of the denominator. x^4-81. In time you should recognize that it is the difference of two perfect squares. Furthermore, in time or maybe right now you are familiar with factoring the difference of two squares a2−b2=(a+b)(a−b) does that strike a familiar chord?

OpenStudy (photon336):

Yeah you can actually apply what @Need_Homework-Help said. \[(x^{4}-81) = (x^{2}-9)(x^{2}+9)\]

OpenStudy (anonymous):

x4−81=(x2+9)(x2−9) Look at both of those factors and what do you notice again?

OpenStudy (anonymous):

what can you say about x^2-9

OpenStudy (halox156):

Thx photon336

OpenStudy (photon336):

of course, this is for the numerator: you can continue to keep simplifying this

OpenStudy (anonymous):

kewl glad 2 help